If you want to buy a house that cost $100,000 today, what will you need to save each year to buy it in 20 years if houses appreciate at 3% per year (Assume you can earn 8%) I know the answer is $3,947 but I want a step by step solution to this
Answer #1

value of the house in 20 years or future value (FV)

=100000*(1+3%)^20

=180611

what will you need to save each year to buy it in 20 years

=FV/(((1+r)^n-1)/r)

=180611/(((1+8%)^20-1)/8%)

=3947

the above is answer..